OpenStudy (anonymous):

A person is given one push on a tire swing and then allowed to swing freely. On the first swing the person travels a distance of 14 feet. On each successive swing, the person travels 80% of the distance of the previous swing. What is the total distance the person swings?

OpenStudy (lgbasallote):

sounds like sum of infinite geometric sequence

OpenStudy (anonymous):

what does that mean

OpenStudy (lgbasallote):

isnt that your lesson?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so what do you mean "what does that mean"

OpenStudy (anonymous):

sum of infinite geometric sequence

OpenStudy (anonymous):

how do i get the distance the person swung

OpenStudy (anonymous):

start adding \(14+14\times .8+14\times .8^2+14\times .8^3+...\)

OpenStudy (anonymous):

or write \[14\times (1+.8+.8^2+.8^3+.8^4+...)\] \[14\times \sum_{k=0}^{\infty}.8^k\]

OpenStudy (anonymous):

you know how to compute this?